# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 5   Structure of Excited States
Most of the resonance couplings have been extracted from the production of single pseudoscalar mesons. In electroproduction, there are six complex helicity amplitudes, requiring a minimum of 11 independent measurements for a complete model-independent determination of the amplitudes (with the exception of an overall phase that cannot be determined). Measurements of isospin amplitudes require additional measurements. Following this, the complex amplitudes would need to be subjected to analyses of their phase motions to determine resonance masses on the (real) energy axis, or poles in the (complex) energy plane. Fortunately, in the lower mass range, a variety of constraints can be applied to limit the number of unknowns when fitting the cross sections and polarization observables. These include the masses of states quite well known from hadronic processes or from meson photoproduction. Also, the number of possible angular momenta is limited to $l_{\pi}\leq 3$ in the examples discussed in the following. Additional constraints come from the Watson theorem [908], which relates the electromagnetic phases to the hadronic ones, and from dispersion relations, which assume the imaginary parts of the amplitude is given by the resonance contribution, and the real parts are determined through dispersion integrals and additional non-resonant pole terms. Other approaches use unitary isobar models, which parameterize all known resonances and background terms and unitarize the full amplitudes in a K-matrix procedure. In the following, we show results based on both approaches, where additional systematic uncertainties have been derived from the differences in the two procedures.
Figure 5.1: Excited states of the proton that have been studied in electroproduction and photoproduction to determine their resonance transition amplitudes or transition form factors and will be discussed in this review. The states are shown with respect to their assignments to supermultiplets in the quark model with $U(6)\otimes O(3)$ symmetry. The insert shows the s-channel diagram for the resonance transition from the proton to the excited state. The transition is described by two (for spin-1/2 states) or three (for higher spins) electrocoupling transition amplitudes. Graphics from Ref. [909].
@@ 5.1   General remarks
On the other hand, these form factors can also be calculated microscopically from the diagrams shown in the bottom of Fig. 5.2. This follows from coupling a photon to the $qqq\to qqq$ scattering amplitude and reading off the residue at the ($N$, $N^{\ast}$) double pole [692]. Since no assumptions are made, this is also an exact expression but now formulated in terms of quarks. Nevertheless, it still contains all hadronic effects: The timelike vector-meson poles have their origin in the quark-photon vertex, which describes the coupling of the photon to a quark. The meson-baryon intermediate states, on the other hand, can only appear in the last diagram where the photon couples to all three quarks, because only this object can create intermediate meson and baryon poles and the corresponding cuts. Note also that all possible gluon topologies are contained in these diagrams, e.g., gluon exchanges between the quarks are absorbed in the wave functions, gluon dressings of the photon coupling in the quark-photon vertex, and more complicated topologies are absorbed in the two- and three-body kernels. In this way, we arrived at a description of the form factor in terms of three valence quarks, which is also valid at any $Q^{2}$.
Therefore, in practice it is less a matter of principle and rather a matter of efficiency how to tune the slide control between ‘quark effects’ and ‘hadronic effects’. Even though in principle all hadronic effects are contained in the diagrams in Fig. 5.2, it is usually more efficient to work with explicit couplings to mesons and baryons like those shown in Fig. 5.4. Furthermore, models usually simplify the diagrams in Fig. 5.2 to various degrees, e.g. by keeping only the first diagram, or by dropping the vector-meson poles in the quark-photon vertex, or using non-relativistic ansätze for the baryon wave functions, etc. At large $Q^{2}$, where all these effects are no longer relevant, one truly probes the ‘three-quark core’. The fact that the quark struck by the photon still needs to communicate with the other two quarks via gluons then leads to quark counting rules which determine the asymptotic behavior of the form factors [910, 911, 912]. At lower $Q^{2}$, the form factors become more and more sensitive to the multihadron or meson-baryon (MB) components. Finally, the form factors have a smooth transition to the timelike region, which is dominated by the vector-meson poles. In the following we adopt this ‘microscopic’ point of view, combined with the available electroproduction data, to interpret our current knowledge on the spacelike transition form factors and corresponding helicity amplitudes.
@@
Figure 5.4: (color online) The running quark mass from Dyson-Schwinger equations and lattice QCD for different current quark masses and in the chiral limit [913].
@@ 5.2   Comparison with state-of-the-art quark-based approaches
In this section we compare the results on electroproduction of several low-mass states with different microscopic approaches using quark degrees of freedom, including light-cone sum rules (LCSR) [914], the Dyson-Schwinger (DSE) approach  [782, 692], the covariant spectator model [915, 18], and lattice QCD. Our baseline for comparisons will be the light-front relativistic quark model (LF RQM) from Refs. [916, 319]. Where available we also compare with models for dynamically generated resonances [917]. Further details and discussions can also be found in the comprehensive review [18].
The LF RQM ingredients, shown in Fig. 5.4, are tuned to the elastic electromagnetic form factors by combining the three-quark and $|\pi N\rangle$ contributions to the nucleon wave function with the form
$|N\rangle=0.95\,|3q\rangle+0.313\,|\pi N\rangle.$ (5.1)
The portions of the three-quark and pion contributions were determined in a light-front cloudy bag model in the fit to the elastic electromagnetic form factors [918]; the condition $F_{1}{}_{p}(0)=1$ provides the normalization.
The electroexcitation of resonances on the proton was investigated in the LF RQM [916, 919] for the resonances of the multiplets $(\textbf{56},0^{+}_{0})$, $(\textbf{56},0^{+}_{2})$, and $(\textbf{70},1^{-}_{1})$. In Refs. [319, 314, 298, 299] the parameters of the LF RQM were specified via description of the nucleon electromagnetic form factors up to $Q^{2}$ = 16 GeV2 by combining the three-quark and pion-cloud contributions in the LF dynamics in Eq. (5.1) and employing a running quark mass as a function of $Q^{2}$ [298, 286]. The inclusion of a running quark mass in the LF RQM is motivated by calculations of the quark mass function using the quark DSE [913] and lattice QCD [742] as displayed in Fig. 5.4. For the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1440)\nicefrac{{1}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, $N(1535)\nicefrac{{1}}{{2}}^{-}$, the weights of the $|3q\rangle$ component in the expansion $|N^{*}\rangle=c_{N^{*}}|3q\rangle+...$ were found from experimental data on $\gamma^{*}p\to\pi^{+}n$ assuming that at $Q^{2}>3$ GeV2 these resonance transitions are determined by $|3q\rangle$ only, while at lower $Q^{2}$ contributions from meson-baryon terms become significant. This has been especially verified in the transitions $\gamma^{\ast}p\to N(1535)\nicefrac{{1}}{{2}}^{-}$ and $\gamma^{\ast}p\to N(1440)\nicefrac{{1}}{{2}}^{+}$, where in both cases several different model calculations coincide in their $Q^{2}$ dependence.
Figure 5.5: Left: Magnetic dipole form factor $G_{M}^{*}(Q^{2})$ of the $\gamma^{*}p\to\Delta(1232)$ transition (here shown in the Ash convention [920] and normalized to a dipole). Center: The ratios of the electric ($R_{EM}$) and scalar ($R_{SM}$) to magnetic multipole transition amplitudes. The data are from Refs. [342, 332, 489, 223, 921] and the solid line is the prediction of the LF RQM [319]. Right: DSE results for the $R_{EM}$ ratio; the full result is compared to the one obtained by including only $s$ waves (or $s$ and $p$ waves) in the nucleon and $\Delta$ wave functions [787].
@@ 5.2.1   The $\gamma^{*}N\to\Delta(1232)$ transition
Figure 5.6: Samples of response functions extracted from the azimuthal $\phi$-angle dependence of the differential cross section in the $\gamma^{*}p\to p\pi^{0}$ channel in the region of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ resonance at different values of $Q^{2}$, from left to right: $Q^{2}=0.4,1.45,3.0,5.0$ GeV2 [15]. The solid curve corresponds to the analysis of the data using the dispersion relation approach. The dashed curves uses the unitary isobar model approach, and the dotted curve is the prediction of the MAID2007 model [382].
Figure 5.7: The $\gamma^{\ast}p\to N(1440)$ helicity transition amplitudes $A_{1/2}$ and $S_{1/2}$ in units of $10^{-3}$ GeV-1/2. Helicity amplitudes at low $Q^{2}$; the data are from  [317, 320, 382]. The open red circle at $Q^{2}\approx 0.1$ GeV2 is the result of an analysis of $ep\to ep\pi^{0}$ data from MAMI [380]. The data are compared to EFT calculations [922], where the solid (black) lines are full results, the long-dashed (red) lines tree contributions and the short-dashed (blue) lines are the loop terms. The dashed-dotted (green) lines are from an empirical parametrization.
As an example, we first discuss the lowest-mass and best studied of all states, the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$. It is most prominently seen both in $\pi N\to\pi N$ elastic scattering and in photoproduction $\gamma N\to N\pi$. The electromagnetic transition of the ground state proton to the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ has been experimentally studied in great detail and is known in a large range of $Q^{2}$ [332, 342, 223]. Figure 5.7 shows samples of the response functions in the mass range of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$.
@@
A well-known shortcoming of quark models and quark-based approaches is that at the real photon point ($Q^{2}=0$), the quark contributions do not account for the full magnitude of the transition amplitudes but only to 65 …75% depending on the approach, as seen in Fig. 5.5. The remainder can be attributed to meson-baryon contributions, also referred to as the ‘pion cloud’, which are not modeled in either of the calculations. On the other hand, at $Q^{2}\geq 2\dots 3$ GeV2 the LF RQM with a momentum-dependent constituent quark mass and the DSE results [787, 782, 786] account for nearly the full strength of the experimental data. This illustrates that the meson-baryon contributions occupy the peripheral, large distance space probed at low $Q^{2}$, while at large $Q^{2}$, i.e. short distances, the quark core accounts for the full resonance response. Without a quark core, the transition amplitudes would fall off much faster with $Q^{2}$ compared to a pure three-quark system.
Also interesting is the electric quadrupole ratio, whose experimental value at low $Q^{2}$ is $R_{EM}\approx-0.02.$ There has been a longstanding prediction of asymptotic pQCD that $R_{EM}\to+1$ at $Q^{2}\to\infty$ [912]. However, the quadrupole ratio $R_{EM}$ shows no sign of departing significantly from its value at $Q^{2}=0$, even at the highest $Q^{2}\approx 6.5$ GeV2. As seen in Fig. 5.5, the LF RQM calculation also barely departs from $R_{EM}=0$ and remains near zero at all $Q^{2}>2$ GeV2. The holographic QCD model [923], on the other hand, predicts a zero-crossing of $R_{EM}$ near $Q^{2}=0.6$ GeV2 followed by a steep rise to $R_{EM}\to+100\%$ at $Q^{2}\to\infty$. This would indicate that the negative constant value shown by the data could be due to meson-baryon contributions which are not included in the theoretical models. However, the DSE result shown in the right panel of Fig. 5.5 offers a different explanation [787]: As discussed in Sec. 4.5, the relativistic nucleon and $\Delta$ wave functions have significant $p$-wave components, which are absent in the non-relativistic quark model. If they are neglected, $R_{EM}$ starts out close to zero and sharply rises towards the asymptotic value; if they are included, the result is able to explain the data. This illustrates that relativistic effects for light baryons have observable consequences, and not every discrepancy between model calculations and experimental data is necessarily a pion-cloud effect.
@@
Figure 5.8: Left and center: The $\gamma^{\ast}p\to N(1440)$ helicity transition amplitudes $A_{1/2}$ and $S_{1/2}$ in units of $10^{-3}$ GeV-1/2. $A_{1/2}$ and $S_{1/2}$ over the full $Q^{2}$ range. The solid curves are from a light-front relativistic quark model [319] and the dashed curves are DSE results [781]. Data are from the CLAS collaboration [317, 320]. Right: Transition multipole amplitude $M_{1-}$ in the $N(1440)\nicefrac{{1}}{{2}}^{+}$ mass region for two values $Q^{2}=0$ and $Q^{2}=2$ GeV2. One can see the dramatic switch from Im $M_{1-}>0$ (corresponding to $A_{1/2}<0$) at the photon point to a large negative Im $M_{1-}$ amplitude at $Q^{2}=2$ GeV2. The solid lines relate to the unitary isobar model and the dashed lines to the fixed-t dispersion relation approach [916] and Inna Aznauryan, private communications.
@@ 5.2.2   The $\gamma^{*}p\to N(1440)\nicefrac{{1}}{{2}}^{+}$ transition and the Roper resonance
The Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$, discovered in 1964 [924] in a phase shift analysis of elastic $\pi N$ scattering data, has been differently interpreted for half a century. A detailed review of the history and current status of the this resonance is presented in a colloquium-style article published in Review of Modern Physics [546]. Earlier model descriptions, such as the Isgur-Karl model which describes the nucleon as a system of three constituent quarks in a confining potential and a one-gluon exchange contribution leading to a magnetic hyperfine splitting of states [535, 534], and the relativized version of Capstick [548], have popularized the models that became the basis for many further developments and variations, e.g. the light front relativistic quark model and the hypercentral quark model [525]. Other models were developed in parallel. The cloudy bag model [925] describes the nucleon as a bag of three constituent quarks surrounded by a cloud of pions. It has been mostly applied to nucleon resonance excitations in real photoproduction, $Q^{2}=0$ [926, 925], with some success in the description of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ and the Roper resonance transitions.
@@
These discrepancies resulted in different interpretations of the Roper resonance. In Ref. [927], a coupled-channel meson exchange model for pion-nucleon scattering was derived from a chirally symmetric Lagrangian, and the $\pi N$ phase shifts and inelasticities were fitted. No genuine three-quark resonance was needed, and the Roper resonance was described by meson-baryon dynamics alone. These findings were confirmed in Ref. [457] where the Roper resonance was listed as a “dynamically generated resonance". In subsequent publications [928, 651], the Roper resonance was decomposed into a large meson-baryon (MB) and a small $qqq$ component.
@@
Is the $N(1440)\nicefrac{{1}}{{2}}^{+}$ a dynamically generated resonance or a three-quark baryon with a MB cloud? This question can be resolved with electroproduction data from CLAS at Jefferson Lab and the parallel development of theoretical approaches. The internal structure of the Roper resonance is revealed by measurements of the $\gamma^{\ast}p\to N(1440)\nicefrac{{1}}{{2}}^{+}$ helicity amplitudes $A_{1/2}$ and $S_{1/2}$. The CLAS collaboration reported these amplitudes over a wide range in $Q^{2}\sim 0.25\dots 4.2$ GeV2 [317, 320], see Fig. 5.8. New electroproduction data on the Roper yield an additional value at very small $Q^{2}$ in the amplitude $S_{1/2}$ shown in the left panel of Fig. 5.8 [380]. Instead of the helicity amplitudes, which satisfy kinematic constraints at timelike values $Q^{2}<0$, it is sometimes more convenient to parametrize the data in terms of two transition form factors $F_{1}$ and $F_{2}$ which are free of kinematic constraints, see e.g. Refs. [490, 929] for explicit relations.
@@
The striking feature visible in the data is the rapid transition of the transverse amplitude $A_{1/2}$ from a large negative magnitude to a large positive one as a function of $Q^{2}$. Translated to the transition form factors, this implies a zero crossing in $F_{2}$, which is a typical signature of a first radial excitation. Such a behavior is also found in quark-based approaches like the DSE approach [781], the covariant spectator model [930], holographic models [931] and the LF RQM with momentum-dependent quark masses [319]. In the DSE approach, the Roper resonance appears as the first radial excitation of the nucleon with a mass above the $N(1535)\nicefrac{{1}}{{2}}^{-}$, see Fig. 4.14, although this does not exclude MB contributions which may still be important [692]. They are especially visible in the helicity amplitudes: At high $Q^{2}$, the process is dominated by the three-quark structure of the $N(1440)\nicefrac{{1}}{{2}}^{+}$, and the calculated amplitudes agree with the data for the highest $Q^{2}$ values. At low $Q^{2}$ the Roper resonance is no longer well described through its quark structure, which indicates the presence of non-quark core effects. If one (tentatively) interprets the difference between the predicted curves and the data as MB effects, then these contributions must be large below $Q^{2}\sim 1$ GeV2. For the Roper resonance such contributions have been described successfully in dynamical meson-baryon models [932] and in effective field theory [922].
@@
There are also efforts to investigate the Roper resonance using lattice QCD. As discussed in Sec. 4.6, the main difficulty is the fact that the Roper resonance lies above several thresholds, where the coupling to $N\pi\pi$ is especially important [880]. This requires generalizations of the Lüscher formalism, which are under active current development [830]. Such caveats must be kept in mind when interpreting lattice results like those shown in Fig. 5.9 b, where the Roper resonance is a stable bound state only at large pion masses (corresponding to large current-quark masses). The bottom panels in Fig. 5.9 show lattice results for the $\gamma^{\ast}p\to N(1440)$ proton to Roper transition form factors $F_{1}$ and $F_{2}$, which are compatible with the experimental data and also display a zero crossing in $F_{2}$ [933].

### v4

@@ 5  Structure of Excited States
Most of the resonance couplings have been extracted from the production of single pseudoscalar mesons. In electroproduction, there are six complex helicity amplitudes, requiring a minimum of 11 independent measurements for a complete model-independent determination of the amplitudes (with the exception of an overall phase that cannot be determined). Measurements of isospin amplitudes require additional measurements. Following this, the complex amplitudes would need to be subjected to analyses of their phase motions to determine resonance masses on the (real) energy axis, or poles in the (complex) energy plane. Fortunately, in the lower mass range, a variety of constraints can be applied to limit the number of unknowns when fitting the cross sections and polarization observables. These include the masses of states quite well known from hadronic processes or from meson photoproduction. Also, the number of possible angular momenta is limited to $l_{\pi}\leq 3$ in the examples discussed in the following. Additional constraints come from the Watson theorem Watson:1954uc , which relates the electromagnetic phases to the hadronic ones, and from dispersion relations, which assume the imaginary parts of the amplitude is given by the resonance contribution, and the real parts are determined through dispersion integrals and additional non-resonant pole terms. Other approaches use unitary isobar models, which parameterize all known resonances and background terms and unitarize the full amplitudes in a K-matrix procedure. In the following, we show results based on both approaches, where additional systematic uncertainties have been derived from the differences in the two procedures.
Figure 5.1: Excited states of the proton that have been studied in electroproduction and photoproduction to determine their resonance transition amplitudes or transition form factors and will be discussed in this review. The states are shown with respect to their assignments to supermultiplets in the quark model with $U(6)\otimes O(3)$ symmetry. The insert shows the s-channel diagram for the resonance transition from the proton to the excited state. The transition is described by two (for spin-1/2 states) or three (for higher spins) electrocoupling transition amplitudes. Graphics from Ref. Burkert:2019kxy .
@@ 5.1  General remarks
On the other hand, these form factors can also be calculated microscopically from the diagrams shown in the bottom of Fig. 5.2. This follows from coupling a photon to the $qqq\to qqq$ scattering amplitude and reading off the residue at the ($N$, $N^{\ast}$) double pole Eichmann:2016yit . Since no assumptions are made, this is also an exact expression but now formulated in terms of quarks. Nevertheless, it still contains all hadronic effects: The timelike vector-meson poles have their origin in the quark-photon vertex, which describes the coupling of the photon to a quark. The meson-baryon intermediate states, on the other hand, can only appear in the last diagram where the photon couples to all three quarks, because only this object can create intermediate meson and baryon poles and the corresponding cuts. Note also that all possible gluon topologies are contained in these diagrams, e.g., gluon exchanges between the quarks are absorbed in the wave functions, gluon dressings of the photon coupling in the quark-photon vertex, and more complicated topologies are absorbed in the two- and three-body kernels. In this way, we arrived at a description of the form factor in terms of three valence quarks, which is also valid at any $Q^{2}$.
Therefore, in practice it is less a matter of principle and rather a matter of efficiency how to tune the slide control between ‘quark effects’ and ‘hadronic effects’. Even though in principle all hadronic effects are contained in the diagrams in Fig. 5.2, it is usually more efficient to work with explicit couplings to mesons and baryons like those shown in Fig. 5.4. Furthermore, models usually simplify the diagrams in Fig. 5.2 to various degrees, e.g. by keeping only the first diagram, or by dropping the vector-meson poles in the quark-photon vertex, or using non-relativistic ansätze for the baryon wave functions, etc. At large $Q^{2}$, where all these effects are no longer relevant, one truly probes the ‘three-quark core’. The fact that the quark struck by the photon still needs to communicate with the other two quarks via gluons then leads to quark counting rules which determine the asymptotic behavior of the form factors Brodsky:1973kr , Brodsky:1974vy , Carlson:1985mm . At lower $Q^{2}$, the form factors become more and more sensitive to the multihadron or meson-baryon (MB) components. Finally, the form factors have a smooth transition to the timelike region, which is dominated by the vector-meson poles. In the following we adopt this ‘microscopic’ point of view, combined with the available electroproduction data, to interpret our current knowledge on the spacelike transition form factors and corresponding helicity amplitudes.
@@
Figure 5.4: (color online) The running quark mass from Dyson-Schwinger equations and lattice QCD for different current quark masses and in the chiral limit Bhagwat:2003vw .
@@ 5.2  Comparison with state-of-the-art quark-based approaches
In this section we compare the results on electroproduction of several low-mass states with different microscopic approaches using quark degrees of freedom, including light-cone sum rules (LCSR) Anikin:2015ita , the Dyson-Schwinger (DSE) approach  Segovia:2016zyc , Eichmann:2016yit , the covariant spectator model Ramalho:2011ae , Ramalho:2023hqd , and lattice QCD. Our baseline for comparisons will be the light-front relativistic quark model (LF RQM) from Refs. Aznauryan:2007ja , Aznauryan:2012ec . Where available we also compare with models for dynamically generated resonances Jido:2007sm . Further details and discussions can also be found in the comprehensive review Ramalho:2023hqd .
The LF RQM ingredients, shown in Fig. 5.4, are tuned to the elastic electromagnetic form factors by combining the three-quark and $|\pi N\rangle$ contributions to the nucleon wave function of the form
$|N\rangle=A\,|3q\rangle+B\,|\pi N\rangle.$ (5.1)
with $A=0.95$ and $B=0.313$. The portions of the three-quark and pion contributions were determined in a light-front cloudy bag model in the fit to the elastic electromagnetic form factors Miller:2002ig ; the condition $F_{1}{}_{p}(0)=1$ provides the normalization.
[…]
Figure 5.5: Left: Magnetic dipole form factor $G_{M}^{*}(Q^{2})$ of the $\gamma^{*}p\to\Delta(1232)$ transition (here shown in the Ash convention Ash:1967rlw and normalized to a dipole). Center: The ratios of the electric ($R_{EM}$) and scalar ($R_{SM}$) to magnetic multipole transition amplitudes. The data are from Refs. CLAS:2001cbm , CLAS:2006ezq , Kelly:2005jy , Frolov:1998pw , Beck:1999ge and the solid line is the prediction of the LF RQM Aznauryan:2012ec . Right: DSE results for the $R_{EM}$ ratio; the full result is compared to the one obtained by including only $s$ waves (or $s$ and $p$ waves) in the nucleon and $\Delta$ wave functions Eichmann:2011aa .
@@ 5.2.1  The $\gamma^{*}N\to\Delta(1232)$ transition
Figure 5.6: Samples of response functions extracted from the azimuthal $\phi$-angle dependence of the differential cross section in the $\gamma^{*}p\to p\pi^{0}$ channel in the region of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ resonance at different values of $Q^{2}$, from left to right: $Q^{2}=0.4,1.45,3.0,5.0$ GeV2 Aznauryan:2011qj . The solid curve corresponds to the analysis of the data using the dispersion relation approach. The dashed curves uses the unitary isobar model approach, and the dotted curve is the prediction of the MAID2007 model Drechsel:2007if .
Figure 5.7: The $\gamma^{\ast}p\to N(1440)$ helicity transition amplitudes $A_{1/2}$ and $S_{1/2}$ in units of $10^{-3}$ GeV-1/2. Helicity amplitudes at low $Q^{2}$; the data are from  CLAS:2012wxw , CLAS:2009ces , Drechsel:2007if . The open red circle at $Q^{2}\approx 0.1$ GeV2 is the result of an analysis of $ep\to ep\pi^{0}$ data from MAMI Stajner:2017fmh . The data are compared to EFT calculations Bauer:2014cqa , where the solid (black) lines are full results, the long-dashed (red) lines tree contributions and the short-dashed (blue) lines are the loop terms. The dashed-dotted (green) lines are from an empirical parametrization.
As an example, we first discuss the lowest-mass and best studied of all states, the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$. It is most prominently seen both in $\pi N\to\pi N$ elastic scattering and in photoproduction $\gamma N\to N\pi$. The electromagnetic transition of the ground state proton to the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ has been experimentally studied in great detail and is known in a large range of $Q^{2}$ CLAS:2006ezq , CLAS:2001cbm , Frolov:1998pw . Figure 5.7 shows samples of the response functions in the mass range of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$.
@@
A well-known shortcoming of quark models and quark-based approaches is that at the real photon point ($Q^{2}=0$), the quark contributions do not account for the full magnitude of the transition amplitudes but only to 65 …75% depending on the approach, as seen in Fig. 5.5. The remainder can be attributed to meson-baryon contributions, also referred to as the ‘pion cloud’, which are not modeled in either of the calculations. On the other hand, at $Q^{2}\geq 2\dots 3$ GeV2 the LF RQM with a momentum-dependent constituent quark mass and the DSE results Eichmann:2011aa , Segovia:2016zyc , Sanchis-Alepuz:2017mir account for nearly the full strength of the experimental data. This illustrates that the meson-baryon contributions occupy the peripheral, large distance space probed at low $Q^{2}$, while at large $Q^{2}$, i.e. short distances, the quark core accounts for the full resonance response. Without a quark core, the transition amplitudes would fall off much faster with $Q^{2}$ compared to a pure three-quark system.
Also interesting is the electric quadrupole ratio, whose experimental value at low $Q^{2}$ is $R_{EM}\approx-0.02.$ There has been a longstanding prediction of asymptotic pQCD that $R_{EM}\to+1$ at $Q^{2}\to\infty$ Carlson:1985mm . However, the quadrupole ratio $R_{EM}$ shows no sign of departing significantly from its value at $Q^{2}=0$, even at the highest $Q^{2}\approx 6.5$ GeV2. As seen in Fig. 5.5, the LF RQM calculation also barely departs from $R_{EM}=0$ and remains near zero at all $Q^{2}>2$ GeV2. The holographic QCD model Grigoryan:2009pp , on the other hand, predicts a zero-crossing of $R_{EM}$ near $Q^{2}=0.6$ GeV2 followed by a steep rise to $R_{EM}\to+100\%$ at $Q^{2}\to\infty$. This would indicate that the negative constant value shown by the data could be due to meson-baryon contributions which are not included in the theoretical models. However, the DSE result shown in the right panel of Fig. 5.5 offers a different explanation Eichmann:2011aa : As discussed in Sec. 4.5, the relativistic nucleon and $\Delta$ wave functions have significant $p$-wave components, which are absent in the non-relativistic quark model. If they are neglected, $R_{EM}$ starts out close to zero and sharply rises towards the asymptotic value; if they are included, the result is able to explain the data. This illustrates that relativistic effects for light baryons have observable consequences, and not every discrepancy between model calculations and experimental data is necessarily a pion-cloud effect.
@@
Figure 5.8: Left and center: The $\gamma^{\ast}p\to N(1440)$ helicity transition amplitudes $A_{1/2}$ and $S_{1/2}$ in units of $10^{-3}$ GeV-1/2. $A_{1/2}$ and $S_{1/2}$ over the full $Q^{2}$ range. The solid curves are from a light-front relativistic quark model Aznauryan:2012ec and the dashed curves are DSE results Segovia:2015hra . Data are from the CLAS collaboration CLAS:2012wxw , CLAS:2009ces . Right: Transition multipole amplitude $M_{1-}$ in the $N(1440)\nicefrac{{1}}{{2}}^{+}$ mass region for two values $Q^{2}=0$ and $Q^{2}=2$ GeV2. One can see the dramatic switch from Im $M_{1-}>0$ (corresponding to $A_{1/2}<0$) at the photon point to a large negative Im $M_{1-}$ amplitude at $Q^{2}=2$ GeV2. The solid lines relate to the unitary isobar model and the dashed lines to the fixed-t dispersion relation approach Aznauryan:2007ja and Inna Aznauryan, private communications.
@@ 5.2.2  The $\gamma^{*}p\to N(1440)\nicefrac{{1}}{{2}}^{+}$ transition and the Roper resonance
The Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$, discovered in 1964 Roper:1964zza in a phase shift analysis of elastic $\pi N$ scattering data, has been differently interpreted for half a century. A detailed review of the history and current status of the this resonance is presented in a colloquium-style article published in Review of Modern Physics Burkert:2017djo . Earlier model descriptions, such as the Isgur-Karl model which describes the nucleon as a system of three constituent quarks in a confining potential and a one-gluon exchange contribution leading to a magnetic hyperfine splitting of states Isgur:1978xj , Isgur:1977ef , and the relativized version of Capstick Capstick:1986ter , have popularized the models that became the basis for many further developments and variations, e.g. the light front relativistic quark model and the hypercentral quark model Giannini:2003xx . Other models were developed in parallel. The cloudy bag model Thomas:1981vc describes the nucleon as a bag of three constituent quarks surrounded by a cloud of pions. It has been mostly applied to nucleon resonance excitations in real photoproduction, $Q^{2}=0$ Bermuth:1988ms , Thomas:1981vc , with some success in the description of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ and the Roper resonance transitions.
@@
These discrepancies resulted in different interpretations of the Roper resonance. In Ref. Krehl:1999km , a coupled-channel meson exchange model for pion-nucleon scattering was derived from a chirally symmetric Lagrangian, and the $\pi N$ phase shifts and inelasticities were fitted. No genuine three-quark resonance was needed, and the Roper resonance was described by meson-baryon dynamics alone. These findings were confirmed in Ref. Ronchen:2018ury where the Roper resonance was listed as a “dynamically generated resonance". In subsequent publications Sekihara:2021eah , Wang:2023snv , the Roper resonance was decomposed into a large meson-baryon (MB) and a small qqqqqq component.
@@
Is the $N(1440)\nicefrac{{1}}{{2}}^{+}$ a dynamically generated resonance or a three-quark baryon with a MB cloud? This question can be resolved with electroproduction data from CLAS at Jefferson Lab and the parallel development of theoretical approaches. The internal structure of the Roper resonance is revealed by measurements of the $\gamma^{\ast}p\to N(1440)\nicefrac{{1}}{{2}}^{+}$ helicity amplitudes $A_{1/2}$ and $S_{1/2}$. The CLAS collaboration reported these amplitudes over a wide range in $Q^{2}\sim 0.25\dots 4.2$ GeV2 CLAS:2012wxw , CLAS:2009ces , see Fig. 5.8. New electroproduction data on the Roper yield an additional value at very small $Q^{2}$ in the amplitude $S_{1/2}$ shown in the left panel of Fig. 5.8 Stajner:2017fmh . Instead of the helicity amplitudes, which satisfy kinematic constraints at timelike values $Q^{2}<0$, it is sometimes more convenient to parametrize the data in terms of two transition form factors $F_{1}$ and $F_{2}$ which are free of kinematic constraints, see e.g. Refs. Tiator:2011pw , Eichmann:2018ytt for explicit relations.
@@
The striking feature visible in the data is the rapid transition of the transverse amplitude $A_{1/2}$ from a large negative magnitude to a large positive one as a function of $Q^{2}$. Translated to the transition form factors, this implies a zero crossing in $F_{2}$, which is a typical signature of a first radial excitation. Such a behavior is also found in quark-based approaches like the DSE approach Segovia:2015hra , the covariant spectator model Ramalho:2017pyc , holographic models deTeramond:2011qp and the LF RQM with momentum-dependent quark masses Aznauryan:2012ec . In the DSE approach, the Roper resonance appears as the first radial excitation of the nucleon with a mass above the $N(1535)\nicefrac{{1}}{{2}}^{-}$, see Fig. 4.14, although this does not exclude MB contributions which may still be important Eichmann:2016yit . They are especially visible in the helicity amplitudes: At high $Q^{2}$, the process is dominated by the three-quark structure of the $N(1440)\nicefrac{{1}}{{2}}^{+}$, and the calculated amplitudes agree with the data for the highest $Q^{2}$ values. At low $Q^{2}$ the Roper resonance is no longer well described through its quark structure, which indicates the presence of non-quark core effects. If one (tentatively) interprets the difference between the predicted curves and the data as MB effects, then these contributions must be large below $Q^{2}\sim 1$ GeV2. For the Roper resonance such contributions have been described successfully in dynamical meson-baryon models Obukhovsky:2011sc and in effective field theory Bauer:2014cqa .
@@
There are also efforts to investigate the Roper resonance using lattice QCD. As discussed in Sec. 4.6, the main difficulty is the fact that the Roper resonance lies above several thresholds, where the coupling to $N\pi\pi$ is especially important Lang:2016hnn . This requires generalizations of the Lüscher formalism, which are under active current development Hansen:2019nir . Such caveats must be kept in mind when interpreting lattice results like those shown in Fig. 5.9 b, where the Roper resonance is a stable bound state only at large pion masses (corresponding to large current-quark masses). The bottom panels in Fig. 5.9 show lattice results for the $\gamma^{\ast}p\to N(1440)$ proton to Roper transition form factors $F_{1}$ and $F_{2}$, which are compatible with the experimental data and also display a zero crossing in $F_{2}$ Lin:2011da .